Reports 1-1 of 1 Clear search Modify search
MIF (General)
takaaki.yokozawa - 6:02 Friday 06 February 2026 (36300) Print this report
Lock failed 260206
In this morning, the lock trial was failed after the ALS_DARM engaged, I tried to search the reason, about the2 Hz oscillation was quite large in ETMX.
During checking the ETMX control signals, I noticed that the glitch of about 2 Hz detected in MN photo sensor L.
But when I checked start time of this glitch, the 2 Hz glitch suddenly disappeared.
After then, locked loss succeeded without any problem.
What this??
Images attached to this report
Comments to this report:
takaaki.yokozawa - 14:54 Friday 06 February 2026 (36307) Print this report
It started the 02:08:24 2026/02/06 JST, and all MN and IM H and V photo sensor can be seen.
(IFO kept locked. LSC_LOCK =9998)
Images attached to this comment
Search Help
×

Warning

×